Please help. I have two questions.
1. Identify the function represented by y=3/x
2. Solve: 8/t+5=t-3/t+5+1/3.
PLEASE HELP!!!

Answer by Alan3354(30993)

Please help. I have two questions.
1. Identify the function represented by y=3/x
------------------------
It's an inverse function. If x increases, y decreases.
2. Solve: 8/t+5=t-3/t+5+1/3.
Is it 8/(t+5) or (8/t) + 5?? It makes a difference, specify it carefully. Same for the right side.